OpenStudy (yusufm):

does sinx+sin2x = sin3x is it additive

OpenStudy (loser66):

What do you think?

OpenStudy (kky10997):

no its not

OpenStudy (yusufm):

that is what I thought kky10997 thank you.

jhonyy9 (jhonyy9):

so just you need prove it that sinx +sin(x+x) = sin(2x+x) use formula sin(a+b) = ... hope you learned it sure

OpenStudy (yusufm):

im trying to simplify\[2\sin2x-2sinx+2*\sqrt{3} cosx-\sqrt{3}=0\] can someone give me a hint as to where to start?